454 mK/s = _______ K/s?

454 mK/s = _______ K/s?
Answer #1

1mill = 10-3 = 0.001

454 mK/s = 454 x 0.001K/s = 0.454 K/s

so, 454 mK/s = 0.454 K/s


answered by: ROBOT